OpenStudy (anonymous):

describe the graph y=6x^2+12x-9 i thought it was a parabola with axis of symmetry of -1

OpenStudy (anonymous):

6x^2+12x=y+9 6(x^2+2x)=y+9 6(x^2+2x+1-1)=y+9 6(x+1)^2-6=y+9 6(x+1)^2=y+15 a parabola of the form x^2=4ay

OpenStudy (anonymous):

does that mean i was right i am confused

OpenStudy (anonymous):

the options are 1, -1, 6, 9

OpenStudy (anonymous):

for the axis of symmetry

OpenStudy (anonymous):

wait ill see

OpenStudy (anonymous):

x=-12/(2*6) =-12/12 =-1

OpenStudy (anonymous):

so yes

OpenStudy (anonymous):

For a quadratic function in standard form, y = ax2 + bx + c, the axis of symmetry is a vertical line x=-b/2a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

any other question?

OpenStudy (anonymous):

umm i have to describe this function y=|x-4|-7 and i got its a absolute function and vertex of 4,-7

OpenStudy (anonymous):

ya right

OpenStudy (anonymous):

OK i wasn't sure cool thanks a lot

OpenStudy (anonymous):

welcome u are in which class and where???

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

i mean which standard do u study

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

u are in which class 10 th or +2 or grade k12 which one?///

OpenStudy (anonymous):

o haha 10

OpenStudy (anonymous):

sorry